Job: Assistant Store Manager

A large department store is located in two major parts of the country: the Midwest and the Southwestern United States. The store employs 700 assistant managers. The store has organized its operations based upon these two geographical regions. Each region runs independently of the other. The data below are used to assess turnover for each division.

Employee Information for assistant store managers

Midwest Division

Southwestern Division

Number of full-time employees on July 1st

500

200

Number of employees that were terminated this year

100

30

Number of employees that quit/retired this year

30

20

Turnover costs

Est cost per opening in Midwest Division

Est cost per opening in Southwestern Division

Hiring Process Costs

Review & revise job description

$40

$50

Post opening on jobs.com

$1000

$1000

Post opening in classifieds

$1200

$1600

Review resumes

$60

$80

Conduct interviews

$200

$150

Conduct background checks

$100

$120

New hire processing

$100

$100

Training Costs

Orientation

$200

$250

Computer system training for new hires

$200

$300

Departmental training

$200

$150

Separation Costs

Exit interview

$50

$50

Overtime costs

$250

$250

[1] Which division has the higher voluntary turnover rate?

Midwest VTR = __________                         Southwestern VTR = __________

[2] Which division has the higher involuntary turnover rate?

Midwest IVTR = __________                       Southwestern IVTR = __________

[3] Which division has the overall higher turnover rate?

Midwest TR = __________                            Southwestern TR = __________

[4] How much does each division spend on turnover for every opening?

Midwest Turnover cost per opening = __________                          

Southwestern Turnover cost per opening = __________

[5] What was the total turnover cost for each division for the given year?

Midwest Total Turnover cost = __________                         

Southwestern Total Turnover cost = __________

[6] If you were the VP of HR, what would this data indicate to you?
Answer #1

Firstly we understand what voluntary turnover rate is and what involuntary turnover rate is.

Voluntary turnover rate: Voluntary turnover rate is the rate at which employees willingly leave a company within a given period. To calculate voluntary turnover rate, divide the number of voluntary separations by the average number of employees during the period.

Involuntary turnover rate: Involuntary turnover rate is the rate at which the number of employees who left the company involuntarily. (Ex: fired employees, Terminated employees). To calculate involuntary turnover rate, divide the number of involuntary separations by the average number of employees during the period.

By analysing the above data,

S No

Particulars

Midwest

Southwestern

A

Average no of employees during the period

500

200

B

Employees who left involuntarily

100

30

C

Employees who left voluntarily

30

20

D

Total employees left during the period (B+C)

100+30 = 130

30+20 = 50

E

Total no of new openings required (D)

130

50

F

Voluntary turnover rate (C/A*100)

30/500*100 = 6%

20/200*100= 10%

G

Involuntary turnover rate (B/A*100)

100/500*100= 20%

30/200*100= 15%

H

Overall turnover rate (D/A*100)

130/500*100= 26%

50/200*100= 25%

I

Total cost per opening

(just simply add all hiring, training, and separation cost)

3600$

4100$

J

Total turnover cost

(It means total number of new openings required for the period * cost per opening)

(E*I)

130*3600 = 468000$

50*4100 = 205000$

Now with the help of above table we try to answer the questions

[1] Which division has the higher voluntary turnover rate?

           Answer: Southwestern

Midwest VTR = 6%                     Southwestern VTR = 10%

[2] Which division has the higher involuntary turnover rate?

           Answer: Midwest

Midwest IVTR = 20%                  Southwestern IVTR = 15%

[3] Which division has the overall higher turnover rate?

           Answer: Midwest

Midwest TR =   26%                   Southwestern TR = 25%

[4] How much does each division spend on turnover for every opening?

           Answer:

Midwest Turnover cost per opening          = 3600$                       

Southwestern Turnover cost per opening = 4100$

[5] What was the total turnover cost for each division for the given year?

            Answer:

Midwest Total Turnover cost           =     468000$          

            Southwestern Total Turnover cost =     205000$

[6] If you were the VP of HR, what would this data indicate to you?

Answer:

This data indicating me that Southwestern is spending more cost per opening comparing to Midwest. Even though Midwest is spending less cost per opening the voluntary turnover rate is anyway less than the rate of Southwestern.